You are analyzing two proposed capital investments with the following cash flows:

Year Project X Project Y

0

- $20,000 - $20,000

1

11,830 6,370

2

5,460 6,370

3

5,460 6,370

4

1,820 6,370


The cost of capital for both projects is 10 percent.

Calculate the profitability index (PI) for each project. (Do not round discount factors. Round intermediate calculations to 2 decimal places, e.g. 15.25 and final answer to 4 decimal places, e.g. 1.2527.)

The PI for project X is _________________ and the PI for project Y is ______________


Which project, or projects, should be accepted if you have unlimited funds to invest?

If you have unlimited funds you should invest in (project X, both projects, neither project, project Y)


Which project should be accepted if they are mutually exclusive?

If they are mutually exclusive you should invest in select a project (Project X, neither project, Project Y)
Answer #1

a)

Project X:

Profitability index = Present value / initial investment

Present value = 11,830 / (1 + 0.1)1 + 5,460 / (1 + 0.1)2 + 5,460 / (1 + 0.1)3 + 1820 / (1 + 0.1)4

Present value = $20,612.20545

Profitability index = 20,612.20545 / 20,000

Profitability index = 1.0306

Project Y:

Profitability index = Present value / initial investment

Present value = 6,370 / (1 + 0.1)1 + 6,370 / (1 + 0.1)2 + 6,370 / (1 + 0.1)3 + 6,370 / (1 + 0.1)4

Present value = $20,192.04289

Profitability index = 20,192.04289 / 20,000

Profitability index = 1.010

b)

Both projects should be accepted as both have PIs greater than 1

c)

If projects are mutually exclusive, project X should be accepted as it has the higher PI.